can you show me the detailed steps for solving for x in the following equation?
2.15x + 2.55*(1-x)=2.516
I can get as far as:
x=2.516-2.55*(1-x)/2.15
but I am unsure how to complete from there.
Thank you in advance for your help.
Scott

Answer by stanbon(75887)

2.15x + 2.55*(1-x)=2.516
2.15x+2.55-2.55x = 2.516
-0.40x=-0.034
x=0.085
Cheers,
Stan H.